Mir Status Report 3 April 1997

Date: 3 April 1997 / Mission Day 82

Mir Station System/Subsystem Activities
Replace PTAB Unit #2 in Priroda
Clean Mir power connectors x-301 and x-302

Mir Systems

Mr. Blagov reported the Vosdukh (CO2 Removal System) was shutdown due to a lack of thermal control from the VGK system in the core module. Some time ago, the core module's VGK and KOH-1V loops (which were originally independent loops) were re-plumbed to create one loop. This was done to reduce the number of required pump panels (one pump panel can support both loops). Earlier in the week, the pressure of the KOH-1V began dropping. Today, the delta pressure at the pump panel dropped to zero, resulting in a shutdown of the loop. The plans are to separate the loops back to two independent loops and determine if another leak exists and isolate it. Mr. Blagov estimated the on-board supply of LiOH canisters would last 10 days, and stated a 15 day supply is manifested on the Progress, docking next week. Currently, the priority of the loop repairs has not been established.

Note: This VGK loop is a separate loop from the one discussed in earlier reports. At last report, the other VGK loop (in Kvant 2) was being patched with an IFM procedure the ground had sent up, and was nearly complete.

Leak checks are in progress on the Kvant 1 VGK.

Last past news: No leaks were found in the VGK. The crew was asked to reconfigure the loop back to the Core Module. The Crew agreed to do this, but requested further work on leak checking be schedule for tomorrow.

Environmental Parameters (Information from 2nd April)
Parameter Min. Max.
Pressure (mmHg) Base Block 748 751
ppO2 (mmHg) 157 159
ppCO2 (mmHg) 4.7 4.9

Planning for MIR23 crew:
MD 83: IFM of VGK Coolant Loop in Kvant 2;
Disassemble Kristall Solar Array Functional Test Setup

MD 84: Additional Time (if required) for IFM of VGK Coolant Loop in Kvant 2
Wet Cleanup; Condensate Presence Monitoring Onboard; O2 Candle Burning

  • submit to reddit


In Fury Born by David Weber - Baen Books
Kepler Communications - Aether

Join our mailing list

Monthly Archives

About this Entry

This page contains a single entry by Keith Cowing published on April 3, 1997 9:16 PM.

Mir Status Report 2 April 1997 was the previous entry in this blog.

Mir Status Report 8 April 1997 is the next entry in this blog.

Find recent content on the main index or look in the archives to find all content.